# The EMPI Code for Plasma-Induced Effects on Radio Waves I: Non-Magnetized Media and Applications to Fast Radio Bursts

### v1

@@ 2.3 Diffractive Regime
The physical essence of the diffractive regime resides in the computation of the Fresnel-Kirchhoff integral, which is discussed in detail by Narayan et al. (1992). The received amplitude $\psi(X,Y)$ at a certain point $(X,Y)$ on the observer plane is calculated by integrating the contributions of light from every point $(x,y)$ on the screen, as:
@@
where $r_{\rm F}$ is the Fresnel scale of this system [^2] The Fresnel scale definition $r_{\rm F}=\sqrt{\lambda d/2\pi}$ varies across literature, regarding the presence or absence of the $2\pi$ factor in the denominator. While the effective propagation distance $d$ may adopt slightly different interpretations depending on the astrophysical context, it generally represents the effective distance of the system.. The first term in the exponential represents the phase change contributed by the plasma screen, while the second term accounts for the geometric phase associated with the propagation of light. In contrast to the refractive regime, the diffractive regime does not trace individual light rays but employs a statistical description. When the characteristic scale of plasma inhomogeneities is exceedingly small, resolving each bump with discrete grid points would require an impractically high resolution. This constraint necessitates implementation of coarser computational grids with larger cell spacing to resolve macroscopic density variations in the plasma screen. Within each grid cell, the effects of subgrid-scale variations are statistically described.
@@
In this regime, the scattering screen manifests frequency-dependent deflective properties. Light scattered within a critical angular range is detectable at the observer, while those deflected beyond this threshold are geometrically excluded. This behavior is quantified through the deflection radius $R_{\rm deflect}$, defined as the region within which the screen provides sufficient deflection to redirect light toward the observer. The formalism assumes complete collection and integration of all radiation within $R_{\rm deflect}$, while neglecting contributions from exterior regions for the inadequate deflection capability.
@@
The numerical integration comprises two critical steps: (i) determination of the integration domain and (ii) evaluation of the phase contributions for each integration element. Following the thin screen formalism for scattering screen developed by Beniamini et al. (2021), the 3D scattering screen is compressed along the main propagation direction to form the equivalent 2D thin screen. After compression, the screen is represented by its column density. The computational elements are rectangular patches bounded by four adjacent grid points, with the position of each patch represented by the center of the corresponding rectangle. Although the theoretical integration domain is nominally defined by the deflection radius $R_{\rm deflect}$, practical implementation requires more considerations. The effective radii $R_{\rm eff}$, which define the integration region, are determined based on several additional considerations. The finite beaming angle $\theta_{\rm FRB}$ creates a circular illumination zone on the plasma screen with radius: $R_{\rm FRB}=d_{\rm sp}\theta_{\rm FRB}$, where $d_{\rm sp}$ is the source-to-screen distance. In addition, the screen has a physical boundary, $R_{\rm screen}$, beyond which no plasma is present and light propagates directly away without deflection, never reaching the observer. Another critical parameter is $R_{\rm deflect}$, which has been defined. The deflection radius $R_{\rm deflect}$ is frequency-dependent and can be estimated as $R_{\rm deflect}=d_{\rm sp}\theta_{\rm deflect}$, where $\theta_{\rm deflect}$ represents the maximum bending capability of the screen. The deflection angle, $\theta_{\rm deflect}$, is approximated as $N\theta_{\rm diff}$, where $N=l_{\rm Thickness}/l_{\rm\nu}$ is the number of scattering events and $\theta_{\rm diff}\sim\lambda/l_{\nu}$ represents the diffraction angle for a single scattering event. Here, $l_{\rm Thickness}$ denotes the thickness of the plasma screen, and $l_{\nu}$ is the characteristic effective bump size for each frequency. The parameter $l_{\nu}$ should be specified either as a free parameter or provided as additional input. Therefore, $R_{\rm deflect}$ is given by:
@@
This computational method adheres to the conventional approaches of refractive regimes, ensuring systematic recording of light propagation times, phase shifts, and DM values. The key difference is that the screen is compressed, and the iteration occurs only once, implying a 2D thin-screen model.
@@ 2.4 Supplementary Treatments for Both Regimes
While preceding sections detailed regime-specific core computational methodologies, this section presents additional shared processing steps. Although earlier analyses emphasized light-plasma interactions within the scattering screen, vacuum propagation segments crucially contribute to geometric components. For computational efficiency, the non-screen regions are modeled as vacuum domains, intentionally neglecting uniform background plasma effects that can be conveniently incorporated through linear superposition. This hybrid treatment enables comprehensive parameter tracking throughout the full propagation path that combines vacuum and plasma segments.
Each light ray carries critical information along its propagation path, including the total travel time $t_{\rm total}=t_{\rm sp}+t_{\rm plasma}+t_{\rm po}$ (where “s” stands for source, “p” for plasma, and “o” for observer), total phase shift $\phi_{\rm total}=\phi_{\rm sp}+\phi_{\rm plasma}+\phi_{\rm po}$, DM value, and additional parameters related to magnetized plasma (which are beyond the scope of this paper). The contributions of the plasma segment have already been discussed. Vacuum contributions, $t_{\rm sp}$ and $t_{\rm po}$ are computed based on the distances between the source and the incident point on the plasma screen, and between the exit point on the screen and the observer, respectively, divided by the speed of light in vacuum $c$. Besides, the wave number in the vacuum is given by $k=\frac{\omega}{c}$, which leads to the two geometric phase shifts: $\phi_{\rm sp}=\omega t_{\rm sp}$ and $\phi_{\rm po}=\omega t_{\rm po}$.
@@ 2.5 Light Reception Process
The synthesis of received signals occurs through the coherent superposition of light that reaches the observer. This critical process ensures the accurate encoding of astrophysical information in the signal. This section details the method of superposing the light to generate the final observed signal. The process begins with the handling of monochromatic light, after which signals from different frequencies are combined according to the resolution of telescope.
@@ 2.5.1 Monochromatic Signal Processing
The Light Receiver module processes discrete-frequency light generated by the simulator through specialized signal construction routines. Since light of the same frequency can interfere and lead to variations in intensity, meticulous treatment becomes imperative. This necessitates the first step of grouping the light by frequency, followed by the coherent summation of all components of the same frequency.
@@ 2.5.2 Superposition Across Multiple Frequencies
Due to the finite spectral and temporal resolutions of observational instruments, rays with different frequencies and arrival times may be mapped to the same grid cell in the waterfall plot, a visual representation of signal intensity over time and frequency. As a result, the intensity of each grid must be computed by summing contributions from all constituent rays occupying that cell. Owing to the incoherent nature of polychromatic radiation, the Stokes parameters of these contributions within the same grid cell can be combined straightforwardly through linear addition.
For an intrinsic pulse, the observed light curve and spectral profile differ from the original signals due to plasma-induced multipath propagation, which introduces time delays across different trajectories. Thus, the observed signal morphology is determined by two factors: the plasma propagation effects and the intrinsic signal profile.
During the millisecond-scale duration of an FRB, the plasma distribution can be treated as static. Consequently, there is no need to recalculate plasma-induced ray trajectories throughout the burst. Instead, they can be computed once, stored, and then used by the receiver module to synthesize the observed signal as the intrinsic signal changes. The intrinsic intensity evolution, characterized by the original light curve and spectrum $I_{\rm intrinsic}(\nu,t)$, is modified by plasma-induced time delays and phase shifts, producing the new intensity distribution $I_{\rm received}(\nu,t)$. At the receiver, the finite temporal and spectral resolutions of the instrument define discrete time-frequency grids, and all contributions falling within the same grid cell are summed to construct the waterfall plot.
@@ 3.1 Screen Generator
The primary function of the Screen Generator is to create a plasma screen that accurately represents the desired electron density distribution, which is crucial for calculating light propagation through the plasma. This screen is described by discrete grid points generated based on specific requirements. Typically, a base distribution, such as a uniform or Gaussian function, serves as the foundation, onto which perturbations can be applied. Structures at various scales, such as clumps, can be superimposed as well. Alternatively, density values can be directly imported from the results of other simulations. The Screen Generator then processes the files to ensure they are in the appropriate format for input into the Light Propagation Simulator. Figure 3 demonstrates three representative types of 3D density distributions. Panel (a) displays a screen populated with irregularly distributed 3D Gaussian blobs. Each blob is well-described by a 3D Gaussian function with a central density peak, and the overall distribution of blobs exhibits random variations in size and position. However, the number of blobs is insufficient, meaning the distribution lacks statistical properties, which makes it difficult to treat analytically. Panel (b) shows a 2D Gaussian density distribution (with no variation along the x-axis), which can be explicitly defined by analytical function. Panel (c) depicts a scattering screen composed of small, randomly distributed clumps. This screen features a uniform average density with random local fluctuations, emphasizing its statistical nature. The statistical properties of such a distribution can typically be described analytically, as is common in scattering screen scenarios. Each discrete grid point in the plasma screen permits individual adjustment, allowing precise control over local electron density characteristics.
Figure 3: Three-dimensional density distributions of the plasma screen. The color bar indicates plasma density, with transparency incorporated to visualize interior density; lower densities correspond to higher transparency. Panel (a) shows a screen with irregularly distributed 3D Gaussian blobs, sized from 0.5 AU to 2.5 AU in 0.5 AU increments, with 50 blobs per size. Each blob is defined by a Gaussian function with a central density of 20 $\rm cm^{-3}$, fluctuating by $\sim$ 0.1, and randomly positioned with some overlap. Panel (b) shows a 2D Gaussian density distribution that remains constant along the x-axis, with a central density of 50 $\rm cm^{-}{3}$ and a cylinder width of 5 AU. Panel (c) illustrates a scattering screen composed of small, randomly distributed clumps, with an average density of 50 $\rm cm^{-}3$ and random fluctuations of approximately 0.1. The fluctuations occur on the scale of the spatial resolution, corresponding to the grid point spacing of 0.5 AU in this plot.
@@ 3.2.2 Implementation Details
To optimize computational efficiency, several strategies are employed: (a) Theoretically, light should be emitted randomly from the source within the half-opening angle, with each light ray corresponding to an incident point on the screen. However, instead of simulating actual light emission from the source, the implementation uses incident points to define the light. This means the distribution of incident points is generated first, followed by tracing the full light trajectory backward to the source. (b) While incident points should theoretically be selected randomly via a Monte Carlo method, they are instead distributed non-randomly to maximize the use of density distribution information and improve computational efficiency. The density is defined precisely at grid points, and each calculation involves four neighboring points that bound a rectangular region. If two incident points fall within the same rectangle, their calculations become nearly identical, rendering redundant unnecessary computations. To minimize redundancy, only one incident point per rectangle is selected. At the same time, to adequately resolve plasma screen features, the grid resolution must be fully utilized. Each region bounded by four neighboring points must include at least one incident point to reflect local density variations. Consequently, the incident plane (the y-z plane) is naturally divided into discrete patches, with the position of each patch represented by its center coordinates and size corresponding to the resolution. The incident points are uniformly distributed, with each patch acting as a computational unit. (c) The bursts have an intrinsic duration, which is extremely short (on the order of microseconds). During a burst, the density distribution of the screen could be considered as static, and the trajectories do not change. Thus, repeated light emission simulations are unnecessary. Trajectory data is stored, and variations arise solely from the intrinsic lightcurve and spectrum. The Light Receiver module synthesizes the observed signal using this stored trajectory pool, explaining why lightcurve and spectrum processing do not occur in the simulator. (d) While all plasma patches on the screen should ideally be considered, principles are applied to help reduce the patch amount. Similar to the process in the diffractive regime, an effective area is defined as the screen region where plasma exists and light can reach. Only the patches in this area are analyzed. Given the FRB beam half-opening angle $\theta_{\rm FRB}$, the screen region within $R_{\rm FRB}=d_{\rm sp}\tan{\theta_{\rm FRB}}$ is illuminated. The patches in the dark region are removed. Patches without plasma are ignored as well, as no deflection occurs and light cannot reach the observer. In case the screen does not intersect the x-axis or has no deflection at the center, a special patch at (0,0) on the y-z plane is always added. (e) Each patch is assigned a unique ID to track its light propagation path. This enables efficient retrieval of observable light data for signal construction in the Light Receiver module. (f) For the diffractive regime, patches within the effective radius $R_{\rm eff}$ are assumed to reach the observer. Trajectory calculations are unnecessary here, but the predefined trajectories should be followed once to obtain the relevant parameters for signal construction.
@@ 3.3 Light Receiver
The Light Receiver module processes light information provided by the Light Propagation Simulator Module, adhering to the methodology detailed in Section 2.5. The input lightcurves and spectra are defined as discrete functions of time and frequency, representing the intrinsic intensity $I_{\rm intrinsic}(\nu,t)$. For each discrete time-frequency grid segment, the intrinsic intensity is changed by the time delay corresponding to its propagation trajectory, generating the modified intensity distribution $I_{\rm received}(\nu,t)$. By taking into account the temporal and spectral resolution of the telescope, contributions from all rays are integrated to synthesize the received waterfall plot. The intensity of a certain grid on the waterfall plot is calculated by summing all contributions within that grid:
@@
Additionally, for the refractive regime, further processing is applied to determine the non-uniform light distribution on the observer plane caused by plasma effects. The Light Propagation Simulator Module outputs intersection points on the observer plane as 2D coordinates, where each point is defined by two values (e.g., (1, 2)). These points are plotted on a single plane, which is partitioned into small bins. The number of points within each bin is counted and normalized relative to a reference bin unaffected by lensing. This procedure produces the 2D light distribution on the observer plane.

### v2

@@ 2.3 Diffractive Regime
The physical essence of the diffractive regime resides in the computation of the Fresnel–Kirchhoff integral, which is discussed in detail by Narayan et al. (1992). The received amplitude $\psi(X,Y)$ at a certain point $(X,Y)$ on the observer plane is calculated by integrating the contributions of light from every point $(x,y)$ on the screen, as:
@@
where $r_{\rm F}$ is the Fresnel scale of this system [^2] The Fresnel scale definition $r_{\rm F}=\sqrt{\lambda d/2\pi}$ varies across literature, regarding the presence or absence of the $2\pi$ factor in the denominator. While the effective propagation distance $d$ may adopt slightly different interpretations depending on the astrophysical context, it generally represents the effective distance of the system.. The first term in the exponential represents the phase change contributed by the plasma screen, while the second term accounts for the geometric phase associated with the propagation of light. In contrast to the refractive regime, the diffractive regime does not trace individual light rays but employs statistical description. When the characteristic scale of plasma inhomogeneities is exceedingly small, resolving each bump with discrete grid points would require an impractically high resolution. This constraint necessitates the implementation of coarser computational grids with larger cell spacing to resolve macroscopic density variations in the plasma screen. Within each grid cell, the effects of subgrid-scale variations are statistically described.
@@
In this regime, the scattering screen manifests frequency-dependent deflective properties. Light scattered within a critical angular range is detectable at the observer, while those beyond this threshold are geometrically excluded. This behavior is quantified through the deflection radius $R_{\rm deflect}$, which is defined as the region within which the screen provides sufficient deflection to redirect light toward the observer. The formalism assumes complete collection and integration of all radiation within $R_{\rm deflect}$, while neglecting contributions from exterior regions due to the inadequate deflection capability.
@@
The numerical integration comprises two critical steps: (i) determination of the integration domain and (ii) evaluation of the phase contributions for each integration element. Following the thin screen formalism for scattering screen developed by Beniamini et al. (2021), the 3D scattering screen is compressed along the main propagation direction to form the equivalent two-dimensional (2D) thinscreen. After compression, the screen is represented by its column density. The computational elements are rectangular patches bounded by four adjacent grid points, with the position of each patch represented by the center of the corresponding rectangle. Although the theoretical integration domain is nominally defined by the deflection radius $R_{\rm deflect}$, practical implementation requires more considerations. The effective radii $R_{\rm eff}$, which define the integration region, are determined based on several additional conditions. The finite beaming angle $\theta_{\rm FRB}$ creates a circular illumination zone on the plasma screen with radius: $R_{\rm FRB}=d_{\rm sp}\theta_{\rm FRB}$, where $d_{\rm sp}$ is the source-to-screen distance. In addition, the screen has a physical boundary, $R_{\rm screen}$, beyond which no plasma is present and light propagates directly away without deflection, never reaching the observer. Another critical parameter is $R_{\rm deflect}$, which has been defined. The deflection radius $R_{\rm deflect}$ is frequency-dependent and can be estimated as $R_{\rm deflect}=d_{\rm sp}\theta_{\rm deflect}$, where $\theta_{\rm deflect}$ represents the maximum bending capability of the screen. The deflection angle, $\theta_{\rm deflect}$, is approximated as $N\theta_{\rm diff}$, where $N=l_{\rm Thickness}/l_{\rm\nu}$ is the number of scattering events and $\theta_{\rm diff}\sim\lambda/l_{\nu}$ represents the diffraction angle for a single scattering event. Here, $l_{\rm Thickness}$ denotes the thickness of the plasma screen, and $l_{\nu}$ is the characteristic effective bump size for each frequency. The parameter $l_{\nu}$ should be specified either as a free parameter or provided as additional input. Therefore, $R_{\rm deflect}$ is given by:
@@
This computational method adheres to the conventional approaches of refractive regimes, ensuring systematic recording of light propagation times, phase shifts, and DM values. The key difference is that the screen is compressed, and the iteration occurs only once, implying a 2D thin screen model.
@@ 2.4 Supplementary Treatments for Both Regimes
While the preceding sections detailed regime-specific core computational methodologies, this section presents additional shared processing steps. Although earlier analyses emphasized light–plasma interactions within the plasma screen, vacuum propagation segments crucially contribute to geometric components. For computational efficiency, the nonscreen regions are modeled as vacuum domains, intentionally neglecting uniform background plasma effects that can be conveniently incorporated through linear superposition. This hybrid treatment enables comprehensive parameter tracking throughout the full propagation path that combines vacuum and plasma segments.
Each light ray carries critical information along its propagation path, including the total travel time $t_{\rm total}=t_{\rm sp}+t_{\rm plasma}+t_{\rm po}$ (where “s” stands for source, “p” for plasma, and “o” for observer), total phase shift $\phi_{\rm total}=\phi_{\rm sp}+\phi_{\rm plasma}+\phi_{\rm po}$, DM value, and additional parameters related to magnetized plasma (which are beyond the scope of this paper). The contributions of the plasma segment have already been discussed. Vacuum contributions, $t_{\rm sp}$ and $t_{\rm po}$ are computed based on the distances between the source and the incident point on the plasma screen, and between the exit point on the screen and the observer, respectively, divided by the speed of light in vacuum $c$. Besides, the wave number in the vacuum is given by $k=\omega/c$, which leads to the two geometric phase shifts: $\phi_{\rm sp}=\omega t_{\rm sp}$ and $\phi_{\rm po}=\omega t_{\rm po}$.
@@ 2.5 Light Reception Process
The synthesis of received signals occurs through the coherent superposition of light that reaches the observer. This critical process ensures the accurate encoding of astrophysical information in the signal. This section details the method of superposing the light to generate the final observed signal. The process begins with the handling of monochromatic light, after which signals from different frequencies are combined according to the resolution of the telescope.
@@ 2.5.1 Monochromatic Signal Processing
The Light Receiver module processes discrete-frequency light generated by the simulator through specialized signal construction routines. Because light of the same frequency can interfere and cause variations in intensity, meticulous treatment is required. This involves two steps: first grouping the light by frequency, then performing coherent summation of all components sharing the same frequency.
@@ 2.5.2 Superposition Across Multiple Frequencies
Due to the finite spectral and temporal resolutions of observational instruments, rays with different frequencies and arrival times may be mapped to the same grid cell in the waterfall plot, a visual representation of signal intensity over time and frequency. Consequently, the intensity of each grid cell must be computed by summing contributions from all constituent rays within it. Owing to the incoherent nature of polychromatic radiation, the Stokes parameters of these contributions in the same grid cell can be combined straightforwardly through linear addition.
For an intrinsic pulse, the observed lightcurve and spectral profile differ from the original signals due to plasma-induced multipath propagation, which introduces time delays across different trajectories. Thus, the observed signal morphology depends on two factors: plasma propagation effects and the intrinsic signal profile.
During the millisecond-scale duration of an FRB, the plasma distribution can be treated as static. Therefore, plasma-induced ray trajectories need not be recalculated throughout the burst. Instead, they are computed once, stored, and then used by the receiver module to synthesize the observed signal as the intrinsic signal evolves. The intrinsic intensity evolution, characterized by the original lightcurve and spectrum $I_{\rm intrinsic}(\nu,t)$, is modified by plasma-induced time delays and phase shifts, producing the observed intensity distribution $I_{\rm received}(\nu,t)$. At the receiver, the finite temporal and spectral resolutions of the instrument define discrete time-frequency grids, and all contributions within each grid cell are summed to construct the waterfall plot.
@@ 3.1 Screen Generator
The primary function of the screen generator is to create a plasma screen that accurately represents the desired electron density distribution, which is crucial for calculating light propagation through the plasma. This screen is described by discrete grid points generated under specific requirements. Typically, a base distribution, such as a uniform or Gaussian function, serves as the foundation, onto which perturbations can be applied. Structures at various scales, such as clumps, can be superimposed as well. Alternatively, density values can be directly imported from the results of other simulations. The screen generator then formats these data for compatibility with the light propagation simulator.
Figure 3 demonstrates three representative types of 3D density distributions. Panel (a) displays a screen populated with irregularly distributed 3D Gaussian blobs. Each blob is well described by a 3D Gaussian function with a central density peak, and the overall distribution of blobs exhibits random variations in size and position. However, the number of blobs is insufficient, meaning the distribution lacks statistical properties, which makes it difficult to be treated analytically. Panel (b) shows a 2D Gaussian density distribution (with no variation along the $x$-axis), which can be explicitly defined by an analytical function. Panel (c) depicts a scattering screen composed of small, randomly distributed clumps. This screen features a uniform average density with random local fluctuations, emphasizing its statistical nature. The statistical properties of such a distribution can typically be described analytically, as is common in scattering screen scenarios. Each discrete grid point in the plasma screen permits individual adjustment, allowing precise control over local electron density.
Figure 3: 3D density distributions of the plasma screen. The color bar indicates plasma density, with transparency incorporated to visualize interior density; lower densities correspond to higher transparency. Panel (a) shows a screen with irregularly distributed 3D Gaussian blobs, sized from 0.5 au to 2.5 au in 0.5 au increments, with 50 blobs per size. Each blob is defined by a Gaussian function with a central density of 20 $\rm cm^{-3}$, fluctuating by $\sim$ 0.1, and randomly positioned with some overlap. Panel (b) shows a 2D Gaussian density distribution that remains constant along the x-axis, with a central density of 50 $\rm cm^{-}{3}$ and a cylinder width of 5 au. Panel (c) illustrates a scattering screen composed of small, randomly distributed clumps, with an average density of 50 $\rm cm^{-}3$ and random fluctuations of approximately 0.1. The fluctuations occur on the scale of the spatial resolution, corresponding to the grid point spacing of 0.5 au in this plot.
@@ 3.2.2 Implementation Details
To optimize computational efficiency, several strategies are employed: (a) Theoretically, light rays should be emitted randomly from the source within the half-opening angle, each intersecting the plasma screen at a unique incident point. However, instead of simulating physical emission from the source, the implementation uses incident points to define the light. This means the distribution of incident points is generated first, followed by tracing the full light trajectory backward to the source. (b) While Monte Carlo sampling is ideal for random incident point selection, a grid-aligned approach is adopted to improve computational efficiency and leverage known density distribution. Each rectangular region bounded by four neighboring grid points is treated as a computational unit. To avoid redundant calculations only one incident point per rectangular region is selected. If two incident points fall in the same rectangle, their calculations become nearly identical, rendering unnecessary computations redundant. At the same time, to adequately resolve plasma screen features, the grid resolution must be fully utilized. Each computational unit must include at least one incident point to reflect local density variations. Thus, the incident plane (the $y$$z$ plane) is naturally divided into discrete patches, each defined by its center coordinates and size matching the grid resolution. The incident points (centers of patches) are uniformly distributed, with each patch acting as a computational unit. (c) Given the millisecond-scale duration of a burst, the density distribution of the screen could be considered static, and the trajectories do not change during the burst. Thus, repeated light emission simulations are unnecessary. Trajectories are precomputed once and stored. Observed signal variations arise solely from the intrinsic lightcurve and spectrum, modulated by stored plasma-induced delays and phase shifts. The light receiver module synthesizes the observed signal using this stored trajectory pool, explaining why lightcurve and spectrum processing do not occur in the simulator. (d) While all plasma patches on the screen should ideally be considered, principles are applied to help reduce the number of patches. Only plasma-containing regions within the FRB beam half-opening angle $\theta_{\rm FRB}$ are considered. Patches located in nonilluminated regions, outside the radius $R_{\rm FRB}=d_{\rm sp}\tan{\theta_{\rm FRB}}$, or within plasma-free regions are removed. In case the screen does not intersect the $x$-axis or has no deflection at the center, a special patch at (0,0) on the $y$$z$ plane is always added. (e) Each patch is assigned a unique ID to track its light propagation path. This enables efficient retrieval of observable light data for signal construction in the light receiver module. (f) For the diffractive regime, patches within the effective radius $R_{\rm eff}$ are assumed to reach the observer. Although trajectory calculations are unnecessary here, the predefined trajectories must be traced once to obtain the parameters required for signal construction.
@@ 3.3 Light Receiver
The light receiver module processes light propagation data from the light propagation simulator module, following the methodology outlined in Section 2.5. The input lightcurves and spectra are discrete functions defined over time and frequency, representing the intrinsic intensity $I_{\rm intrinsic}(\nu,t)$. For each discrete time-frequency bin, the intrinsic intensity is adjusted by the time delay associated with its propagation trajectory, generating the modified intensity distribution $I_{\rm received}(\nu,t)$. By taking into account the temporal and spectral resolution of the telescope, contributions from all rays are integrated to synthesize the received waterfall plot. The intensity of a certain grid on the waterfall plot is calculated by summing all contributions within that grid:
@@
Additionally, for the refractive regime, further processing is applied to determine the nonuniform light distribution on the observer plane caused by plasma effects. The light propagation simulator module outputs intersection points on the observer plane as 2D coordinates, where each point is defined by two values (e.g., (1, 2)). These points are plotted on a single plane, which is partitioned into small bins. The count of points within each bin is normalized relative to a reference bin unaffected by plasma lensing. This procedure produces the 2D light distribution on the observer plane.